LPIC Exams

• LPIC1 Junior Level Administration – 101 General Linux Part 1

– 102 General Linux Part 2

• LPIC2 Intermediate Level Administration – 201 Advanced Linux Administration – 202 Linux Networking Administration

• LPIC3 Senior Level Administration (Two topics to be completed) – 321 Windows Integration

– 322 Internet Server

– 323 Database Server

– 324 Security, Firewalls and Encryption – 325 Kernel Drivers and Packages

– 32? TBA

LPIC 1–Books

References

[1] Jeffrey Dean. LPI Linux Certification in a Nutshell—A Desktop Quick Reference. O’Reilly, 2001.

[2] Angie Nash and Jason Nash. LPIC 1 Certification Bible. Hungry Minds.

[3] Paul Sheer. LINUX Rute Users Tutorial and Exposition Prentice Hall.

[4] Theresa Hadden Martinez. Prep Kit General Linux 1 Que.

[5] Ross Brunson. Lpic 1 Exam Cram 2: Linux Professional Institute Certification Exams 101 and 102 (Exam Cram 2) Que.

Certifications • LPIC

• RHCT, RHCE

• Novell CLE (see http://www.novell.com/training/certinfo/cle/)

• SAGE LPIC Level 1 Certification

All current LPI exams use computer-based tests administered through VUE or Prometric. Each exam costs $100 (US dollars) and the current outline is for there to be two exams at each level.

• Suitable tasks:

– can work at the Unix command line

– performs easy maintenance tasks: help out users, add users to a larger system, backup & restore, shutdown & reboot

– can install and configure a work station (incl. X) and connect it to the LAN, or a

stand-alone PC via modem to the Internet.

• Note:

– This is somewhat heavier than the SAGE “Novice” level: we aim beyond the power-user and helpdesk level. It is about at the “prerequisites” level of the Red Hat training program.

LPI Level 1 Topics

• LPI 101 – 101: Hardware & Architecture

– 102: Linux Installation & Package Management

– 103: GNU & Unix Commands

– 104: Devices, Linux Filesystems, Filesystem Hierarchy Standard

– 110: The X Window System

• LPI 102 – 105: Kernel

– 106: Boot, Initialisation, Shutdown and Runlevels

– 107: Printing

– 108: Documentation

– 109: Shells, Scripting, Programming and Compiling

– 111: Administrative Tasks

– 112: Networking Fundamentals

– 113: Networking Services

– 114: Security LPI Level 2 Topics

• 2.201: Kernel (0)

• 2.202: System Startup (0)

• 2.203: Filesystem (0)

• 2.204: Hardware (0)

• 2.205: Network (0)

• 2.206: Mail & News (0)

• 2.207: DNS (0)

• 2.208: Web Servers (0)

• 2.209: File & Service Sharing (0)

• 2.210: Network Client Management (0)

• 2.211: System Maintenance (0)

• 2.212: System Security (0)

• 2.213: System Customisation & Automation (0)

• 2.214: Troubleshooting (0)

Example of a Topic–107: Printing There are three Objectives in the Printing topic:

• 1.107.2 Manage printers and print queues

• 1.107.3 Print files

• 1.107.4 Install and configure local and remote printers

Example of an Objective 1.107.2 Manage printers and print queues: • Modified: 2003-Mar-17

• Maintainer: Kara Pritchard

• Weight: 1 Description: Candidates should be able to manage print queues and user print jobs. This

objective includes monitoring print server and user print queues and troubleshooting general printing problems.

Key files, terms, and utilities include: • /etc/printcap

       • lpc
       • lpq
       • lprm
       • lp

License covering this document Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005 Geoffrey Robertson <ge@ffrey.com> and Nick Urbanik <nicku@nicku.org>

You can redistribute modified or unmodified copies of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation — either version 2 of the License or (at your option) any later version.